OpenStudy (anonymous):

On a right triangle if a= 3 and b=3 what would c be?

jimthompson5910 (jim_thompson5910):

using the pythagorean theorem a^2 + b^2 = c^2 3^2 + 3^2 = c^2 9 + 9 = c^2 ... ... ... c = ???

OpenStudy (anonymous):

Okay so whatever c is I have to square that?

jimthompson5910 (jim_thompson5910):

the opposite actually, you have to take the square root

jimthompson5910 (jim_thompson5910):

c^2 = 18 so c = ???

OpenStudy (anonymous):

C would equal 4.24?

jimthompson5910 (jim_thompson5910):

correct Exact answer: \(\large c = 3\sqrt{2}\) Approximate answer: \(\large c \approx 4.24264068711929\) and you can round the approximate answer however you need to

OpenStudy (anonymous):

Oh okay well thank you very much!

jimthompson5910 (jim_thompson5910):

you're welcome